# Two-dimensional fractional Brownian motion: Analysis in time and frequency domains

### v1

@@ VI Conclusions
In this work, we introduced a natural construction of 2D fBm that accounts for anisotropic scaling and cross-dimensional dependencies, extending the classical framework of Mandelbrot and van Ness to multivariate settings. By incorporating correlated Gaussian noises and a matrix-valued Hurst operator, we provided two distinct formulations, causal and well-balanced 2D fBm, and derived their respective auto- and cross-covariance across the two components. Our analysis highlights key differences between these two formulations, particularly in the asymmetry of causal fBm and implications for time-reversal. Furthermore, we analyzed the spectral properties of the processes, presenting both full analytical derivations and asymptotic behaviors in the frequency domain. Our results reveal how variations in the Hurst exponents and correlation structure manifest in the spectral content, offering new insights into the interpretation of multivariate time series in diverse systems. An especially interesting case arises when $H_{1}=H_{2}$. In this situation, the process remains a 2D fBm under arbitrary rotations. In contrast, when $H_{1}\neq H_{2}$, a rotation mixes the coordinates in such a way that the resulting marginals are no longer fBm, but instead they exhibit more complex dependencies.
Extensive numerical simulations validated our theoretical findings, demonstrating consistency in both the time and frequency domains. The proposed framework thus provides a robust tool for modeling and analyzing complex anomalous diffusion processes in multidimensional environments, especially in settings where directional dependence and inter-component correlations cannot be ignored. This study opens several avenues for further exploration. A potential immediate extension involves generalizing the proposed constructions to higher dimensions. Our work opens the way to the application of 2D fBm to empirical data in fields such as neuroscience, single-particle tracking, structural dynamics, and financial econometrics, which may yield new insights into underlying spatiotemporal correlations. While all analyses are shown for the two-dimensional case, they can be readily generalized to any dimension.
@@ Appendix A Proof of Theorem 1, causal case and some additional remarks
Utilizing (9.3.28-29) from [42], we can calculate $A=A_{1}+iA_{2}$ that later will be used for determining the covariance structure of the proposed process
@@
$A_{2}$ $=\begin{bmatrix}\Gamma(H_{1}+\frac{1}{2})\cos\left(\frac{(H_{1}-\frac{1}{2})\pi}{2}\right)\sigma_{1}a_{H_{1}}&0\\ \rho\Gamma(H_{2}+\frac{1}{2})\cos\left(\frac{(H_{2}-\frac{1}{2})\pi}{2}\right)\sigma_{2}a_{H_{2}}&\sqrt{1-\rho^{2}}\Gamma(H_{2}+\frac{1}{2})\cos\left(\frac{(H_{2}-\frac{1}{2})\pi}{2}\right)\sigma_{2}a_{H_{2}}\end{bmatrix}$. (57)
@@
$A$ $=\begin{bmatrix}\Gamma(H_{1}+\frac{1}{2})\exp\{i\frac{\pi}{2}(\frac{3}{2}-H_{1})\}\sigma_{1}a_{H_{1}}&0\\ \rho\Gamma(H_{2}+\frac{1}{2})\exp\{i\frac{\pi}{2}(\frac{3}{2}-H_{2})\}\sigma_{2}a_{H_{2}}&\sqrt{1-\rho^{2}}\Gamma(H_{2}+\frac{1}{2})\exp\{i\frac{\pi}{2}(\frac{3}{2}-H_{2})\}\sigma_{2}a_{H_{2}}\end{bmatrix}$. (58)
@@
The connection between the elements of matrix $C$ given in (59) and the parameters $\rho_{jk}$ and $\eta_{jk}$ are as follows
@@ Appendix B Construction of a well-balanced 2D fBm
Defintition 3 (Well-balanced 1D fBm).
@@ Appendix D Proof of Theorem 3, spectral content of the increment process of 2D fBm
Remembering that $AA^{*}=C$ (elements of $C$ are given in (59) and (82)), and taking Remark 1.3.6 and 1.3.8 [42] we obtain the final result for the power spectral density of $\Delta\mathbf{Z}$.
@@ Appendix E Proof of Theorem 4, spectral content of 2D fBm trajectory
$\mathbb{III}$ $\equiv\int_{0}^{1}\int_{0}^{1}e^{i\tilde{\omega}(x-y)}w_{jk}(x-y)|x-y|^{H_{j}+H_{k}}\mathrm{d}x\,\mathrm{d}y=\quad|\textrm{substitute }u=x-y,\ z=x|$
@@ Appendix F Proof of Theorem 5, asymptotics of ensemble-averaged PSD
Thus, grouping the elements by their powers, we have (as $\tilde{\omega}\to 0$)

### v2

@@ VI Conclusions
In this work, we introduced a natural construction of 2D fBm that accounts for anisotropic scaling and cross-dimensional dependencies, extending the classical framework of Mandelbrot and van Ness to multivariate settings. By incorporating correlated Gaussian noises and a matrix-valued Hurst operator, we provided two distinct formulations, causal and well-balanced 2D fBm, and derived their respective auto- and cross-covariance across the two components. Our analysis highlights key differences between these two formulations, particularly in the asymmetry of causal fBm and implications for time-reversal. Furthermore, we analyzed the spectral properties of the processes, presenting both full analytical derivations and asymptotic behaviors in the frequency domain. Our results reveal how variations in the Hurst exponents and correlation structure manifest in the spectral content, offering new insights into the interpretation of multivariate time series in diverse systems. The results also clarify the qualitative distinction between the causal and well-balanced versions of the 2D fBm. While both share the same marginal scaling properties and stationary increments, their dependence structures differ substantially when $H_{1}\neq H_{2}$. The causal model exhibits asymmetric cross-covariance structure and a phase shift in the spectral domain. In contrast, the well-balanced construction remains time-reversible for all parameter values, with symmetric cross-covariance and no cross-PSD structure. In this sense, the two formulations are complementary. The difference between the models provides a means to identify a model in a real data set. A notable case arises when $H_{1}=H_{2}$. In this situation, the process remains a 2D fBm under arbitrary rotations. In contrast, when $H_{1}\neq H_{2}$, a rotation mixes the coordinates in such a way that the resulting marginals are no longer fBm, but instead they exhibit more complex dependencies. A 2D processes with $H_{1}=H_{2}$ can thus describe a system with two-component fBm in the same geometric space, such as the two-dimensional motion of particles in a complex environment with underlying non-uniform structure. On the other hand, $H_{1}\neq H_{2}$ describes a more general case where each component may represent a different fBm, albeit exhibiting correlations. Such two-component processes can emerge in diverse systems. To name a few examples: in climate science, they can describe temperature [67, 68] and precipitation [10]; in hydrology, river flows [10, 69] and erosion [70]; and in physiology, heart rate [71] and breathing patterns [72].
Extensive numerical simulations validated our theoretical findings, demonstrating consistency in both the time and frequency domains. The proposed framework thus provides a robust tool for modeling and analyzing complex anomalous diffusion processes in multidimensional environments, especially in settings where directional dependence and inter-component correlations cannot be ignored. This study opens several avenues for further exploration. A potential immediate extension involves generalizing the proposed constructions and results to higher dimensions. Our work opens the way to the application of 2D fBm to empirical data in fields such as neuroscience, single-particle tracking, structural dynamics, and financial econometrics, which may yield new insights into underlying spatiotemporal correlations.
@@ Appendix A Proof of Theorem 1, causal case and some additional remarks
Utilizing (9.3.28-29) from [52]i.e.,
$M_{+}-M_{-}$ $=\sqrt{2\pi}\left[\sin\left(\frac{D\pi}{2}\right)\right]^{-1}\Gamma(D+I)^{-1}A_{1},$ (56)
$M_{+}+M_{-}$ $=\sqrt{2\pi}\left[\sin\left(\frac{D\pi}{2}\right)\right]^{-1}\Gamma(D+I)^{-1}A_{2},$ (57)
we can calculate $A=A_{1}+iA_{2}$ that later will be used for determining the covariance structure of the proposed process
@@
$A_{2}$ $=\frac{1}{\sqrt{2\pi}}\begin{bmatrix}\Gamma(H_{1}+\frac{1}{2})\cos\left(\frac{(H_{1}-\frac{1}{2})\pi}{2}\right)\sigma_{1}a_{H_{1}}&0\\ \rho\Gamma(H_{2}+\frac{1}{2})\cos\left(\frac{(H_{2}-\frac{1}{2})\pi}{2}\right)\sigma_{2}a_{H_{2}}&\sqrt{1-\rho^{2}}\Gamma(H_{2}+\frac{1}{2})\cos\left(\frac{(H_{2}-\frac{1}{2})\pi}{2}\right)\sigma_{2}a_{H_{2}}\end{bmatrix}$. (61)
@@
$A$ $=\frac{1}{\sqrt{2\pi}}\begin{bmatrix}\Gamma(H_{1}+\frac{1}{2})\exp\{i\frac{\pi}{2}(\frac{3}{2}-H_{1})\}\sigma_{1}a_{H_{1}}&0\\ \rho\Gamma(H_{2}+\frac{1}{2})\exp\{i\frac{\pi}{2}(\frac{3}{2}-H_{2})\}\sigma_{2}a_{H_{2}}&\sqrt{1-\rho^{2}}\Gamma(H_{2}+\frac{1}{2})\exp\{i\frac{\pi}{2}(\frac{3}{2}-H_{2})\}\sigma_{2}a_{H_{2}}\end{bmatrix}$. (62)
@@
The connection between the elements of matrix $C$ given in Eq. (63) and the parameters $\rho_{jk}$ and $\eta_{jk}$ are as follows
@@ Appendix B Construction of a well-balanced 2D fBm
Definition 3 (Well-balanced 1D fBm).
@@ Appendix D Proof of Theorem 3, spectral content of the increment process of 2D fBm
Remembering that $AA^{*}=C$ (elements of $C$ are given in Eqs. (63) and (86)), and taking Remark 1.3.6 and 1.3.8 [52] we obtain the final result for the power spectral density of $\Delta\mathbf{Z}$.
@@ Appendix E Proof of Theorem 4, spectral content of 2D fBm trajectory
$\mathbb{III}$ $\equiv\int_{0}^{1}\int_{0}^{1}e^{i\tilde{\omega}(x-y)}w_{jk}(x-y)|x-y|^{H_{j}+H_{k}}\mathrm{d}x\,\mathrm{d}y=\qquad|\textrm{substitute }u=x-y,\ z=x|$
@@ Appendix F Proof of Theorem 5, asymptotics of ensemble-averaged PSD
where $a_{jk}=\Gamma(H_{j}+H_{k})\sin\left(\frac{\pi}{2}(H_{j}+H_{k})\right)$.
Thus, grouping the elements by their powers, we have (as $\tilde{\omega}\to\infty$)